OpenStudy (anonymous):

If 2x+3 is an integer, write a simplified expression to represent the sum of the next three consecutive integers to 2x+3.

OpenStudy (jack1):

y = 2x +3 y is an integer so 3 consecutive terms are if x = a+0, a+1 and a+2 y = 2x +3 y = 2(a+0) + 3 = 2a + 3 y = 2(a+1) + 3 = 2a + 2 +3 y = 2(a+2) + 3 = 2a + 4 +3 so equation for consecuative intergers = y = 2x + 2^n + 3 where n is the number in the sequence and x is the initial number in the series

OpenStudy (jack1):

n = number in the sequence above or below our nominated x value *

OpenStudy (anonymous):

thank you so much

OpenStudy (jack1):

all good, y w dude